OpenStudy (anonymous):

integrate( tsqrt*t + sqrt(t))/t^2

OpenStudy (abb0t):

\[\int\limits \frac{ t \sqrt{t} + \sqrt{t} }{ t^2 }dt \] is that correct?

OpenStudy (abb0t):

\[\int\limits \frac{ t \sqrt{t} }{ t^2 } dt + \int\limits \frac{ \sqrt{t} }{ t }dt\]

OpenStudy (anonymous):

tsqrt(t( * t^-2 + sqrt(t) * t^-1?

OpenStudy (anonymous):

oh and yeah thats correct

OpenStudy (abb0t):

\[\int\limits \frac{ \sqrt{t} }{ t } dt + \int\limits \frac{ t^\frac{ 1 }{ 2 } }{ t^2 }dt \]

OpenStudy (abb0t):

now, you can subtract the t's to get a single "t" in which you can easily integrate when you have it in the form: \(\large t^n\)

OpenStudy (anonymous):

subtract ? cant u factor 1/t

OpenStudy (abb0t):

remember \(\sqrt{t} = t^\frac{1}{2}\)

OpenStudy (abb0t):

No, t is not a constant in this case, so it cannot be factored out.

OpenStudy (anonymous):

3t^(3/2)/2t for the first part

OpenStudy (abb0t):

\[\int\limits (t^{\frac{ 1 }{ 2 }-(-1)} dt + \int\limits t^{\frac{1}{2}-(-2)}dt\]

OpenStudy (anonymous):

how did u subtract t

OpenStudy (abb0t):

because remember that anything thats int he denominator is negative. So if for example I had \(\large \frac{1}{t}\) what that means is basically \(\large t^{-1}\)

OpenStudy (anonymous):

oh that what you mean

OpenStudy (abb0t):

and if you remember from algebra and the powers, if you have \(\huge \frac{t^3}{t^2}\) it's same thing as \(t^{3-2} = t\)

OpenStudy (anonymous):

t^3/2 + t^5/2

OpenStudy (anonymous):

t^5/2/5/2 + t^7/2/7/2

OpenStudy (abb0t):

You're almost right, but you're missing a constant.

OpenStudy (anonymous):

+ c

OpenStudy (abb0t):

since this is ain indefinite integral, you need to have a constant at the end.

OpenStudy (anonymous):

u mean + c right

OpenStudy (anonymous):

tyvm

OpenStudy (abb0t):

Yeah, it shoudl always be the function + constant. f(x) + C but sine youre workiong with "t" as a variable, its f(t) + C

OpenStudy (anonymous):

k tyty

OpenStudy (anonymous):

hop i ace this quiz tmw